ID |
Date |
Author |
Type |
Category |
Subject |
3294
|
Mon Jul 26 20:12:18 2010 |
kiwamu | Update | SUS | OSEMs on PRM |
[Alberto and Kiwamu]
We installed the OSEMs to the new PRM.
As I wrote down on the elog (see here) today's mission was to install the OSEMs to the PRM.
After putting them on the tower we adjusted the readout offsets by sliding the OSEMs so that they can stay in the linear sensing ranges.
Now all of the OSEMs have almost good separation distances from the PRM.
In the attached picture you can see the OSEMs installed on the PRM tower ( middle: PRM tower, left: BS tower)
(what we did)
1. moved the PRM tower close to the door so that we could easily access the PRM.
2. leveled the table by putting some weights and confirmed the level by a bubble level tool.
- We must level the table every time when we set / adjust any OSEMs, otherwise the readout voltages of the OSEMs vary every time due to the tilted table.
3. released the PRM by loosing the earthquake stops
4. put the OSEMs with approximately right separation distances from the PRM.
- At this phase we can see the readout of the OSEMs, which were oscillating freely because we still didn't enable the damping.
- The OSEM positions were checked by looking at useful notes on the wiki (see here).
5. turned on the damping servo of the OSEMs
- Without changing any gains, it worked well.
- Then we could see stable readouts of the OSEMs which didn't show any oscillations in turn because of the damping.
6. checked the level of the table again
7. set each of the OSEM readouts to the half of its maximum value by sliding their positions slightly.
- The readout offsets were at almost the half value within +/- 100 mV accuracy (this was the best accuracy we could adjust by our hands)
8. screwed down the earthquake stops to lock the PRM.
- Now the damping is off.
9. closed the door
(to be done)
* Putting the PRM tower back to the designed place
* Installation of the pick off mirror
* Arrangement of the PZT mirror |
Attachment 1: DSC_2279.JPG
|
|
3293
|
Mon Jul 26 14:24:46 2010 |
josephb, kiwamu | Update | CDS | RFM test take 1 |
Kiwamu and I strung a temporary RFM fiber from the c1iscex machine (in the new 1X9 rack) to the c1sus machine (in the new 1X4 rack). This was connected into the respective RFM cards. Once we put the fiber in correctly, the status lights came on the RFM card, which is a good sign. This did not go through the RFM bypass, and did not interfere with any other RFM connections.
We created a simple model to test the RFM card, which basically was 4 RFM memory locations passing back and forth between 2 filters on each machine. These models were called c1rf0 (on c1sus) and c1rf1 (on c1iscex). We added 4 entries to the /cvs/cds/caltech/chans/ipc/C1.ipc file corresponding to the 4 RFM memory locations, set their ipcType=RFM and set the ipcRate to 65536. The ipcNum were set from 0 to 3. The models ran, however, the data we were trying to pass over the RFM card did not seem to be being passed. Currently trying to contact Alex via e-mail to get debugging advice, and confirm the ipc file is setup correctly. |
3292
|
Mon Jul 26 12:31:36 2010 |
kiwamu | Update | CDS | front end machine for the X end |
A brief report about the new front end machine "C1ISCEX" installed on the X end (old Y end).
Still the DAC is not working.
- Timing card
It's working correctly.
The 1PPS clock signal is supplied by the vertex clock distributer via a 40m long fiber.
- ADC
All 16 channels are working well.
We can see the signals in the medm screen while injecting some signals to the ADC by using a function generator.
-DAC
All 16 channels do NOT work.
We can not see any signals at the DAC SCSI cable while digitally injecting a signal on the medm screen. |
3291
|
Mon Jul 26 11:15:23 2010 |
Gopal | HowTo | COMSOL Tips | Pictures from Transfer Function Tutorial on the Wiki |
The attached pictures give a brief overview of my transfer function measurement procedure in COMSOL. For more details, please see the Wiki.







|
3290
|
Mon Jul 26 10:04:24 2010 |
steve | Update | General | Initial Crane Inspection reveals flaws: wiring, oil and imbalance |
Quote: |
The guy from KroneCrane (sp?) came today and started the crane inspection on the X End Crane. There were issues with our crane so he's going to resume on Monday. We turned off the MOPA fur the duration of the inspection.
- None of our cranes have oil in the gearbox and it seems that they never did since they have never been maintained. Sloppy installation job. The crane oiling guy is going to come in on Monday.
- They tried to test the X-End crane with 2500 lbs. (its a 1 ton crane). This tripped the thermal overload on the crane as intended with this test. Unfortunately, the thermal overload switch disabled the 'goes down' circuit instead of the 'goes up' circuit as it should. We double checked the wiring diagram to confirm our hypothesis. Seems the X-End crane was wired up incorrectly in the first place 16 years ago. We'll have to get this fixed.
The plan is that they will bring enough weight to test it at slightly over the rating (1 Ton + 10 %) and we'll retry the certification after the oiling on Monday.
|
The south end crane has one more flaw. The wall cantilever is imbalanced: meaning it wants to rotate south ward, because its axis is off.
This effects the rope winding on the drum as it is shown on Atm2
Atm1 is showing Jay Swar of KoneCrane and the two 1250 lbs load that was used for the test. Overloading the crane at 125% is general practice at load testing.
It was good to see that the load brakes were working well at 2500 lbs. Finally we found a good service company! and thanks for Rana and Alberto
for coming in on Saturday. |
Attachment 1: 2500.JPG
|
|
Attachment 2: sedrum.JPG
|
|
3289
|
Mon Jul 26 10:02:42 2010 |
josephb | Update | CDS | Rerouted RFM around c1lsc, took RFM card out of c1lsc |
If you're refering to just the medm screen, those can be restored from the SVN. As we're moving to a new directory structure, starting with /opt/rtcds/caltech/c1/, the old LSC screens can all be put back in the /cvs/cds/caltech/medm/c1/lsc directory if desired.
The slow lsc aux crate, c1iscaux2, is still working, and those channels are still available. I confirmed that one was still updating. As a quick test, I went to the SVN and pulled out the C1LSC_RFADJUST.adl file, renamed it to C1LSC_RFadjust.adl and placed it in /cvs/cds/caltech/medm/c1/lsc/, and checked it linked properly from the C1IOO_ModeCleaner.adl file. I haven't touched the modulation depths, as I didn't want to mess with the mode cleaner, but if I get an OK, we can test that today and confirm that modulation depth control is still working.
Quote: |
I just realized that an unfortunate casualty of this LSC work was the deletion of the slow controls for the LSC which we still use (some sort of AUX processor). For example, the modulation
depth slider for the MC is now in an unknown state.
|
|
3288
|
Mon Jul 26 08:51:31 2010 |
steve | Update | PEM | lowering AC setpoint by 2F makes PSL happy |
Quote: |
The control room temp is warmer than usual. The heat exchanger Office Pro 18 set point was lowered from 70 to 68F yesterday.
The MOPA headtemp is higher also. The Neslab chiller bath temp peaks around 21.6 C daily. This should be rock solid 20.00 C
It did not have any effect.
Now, I have just lowered the thermostat setting of room 101 from 73 to71F I hope Koji can take this.
|
Little change in the AC set temp can make wonders. Neslab chiller bath temp 19.99C is back to normal and daily variation of PSL are much better. |
Attachment 1: htempRTemp.jpg
|
|
3287
|
Sun Jul 25 18:47:23 2010 |
Alberto | Update | SVN | Optickle 40mUpgrade model updated to include short cavity length corrections |
I uploaded an updated optickle model of the upgrade to the SVN directory with the optickle models (here). |
3286
|
Sat Jul 24 14:27:36 2010 |
rana | Update | Electronics | FSS Oscilaltor Phase Noise Measurement |
I reconnected the RF signal to the FSS and to the FSS' EOM so that we could lock the refcav again.
I then started a 3 sec. period trianglewave on the AOM drive amplitude to see if there is a direct coupling from RIN to Frequency. Ideally we will be able to measure this by looking at the RCTRANS and the FSS-FAST. |
3285
|
Sat Jul 24 14:03:19 2010 |
Alberto | Update | Electronics | FSS Oscilaltor Phase Noise Measurement |
[Rana, Alberto]
Today we measured the phase noise of the oscillator used for the FSS.
The source is a Wenzel crystal at about 21.5MHz that Peter Kalmus built some time ago.
We basically used the same technique that Frank and Megan have been using lately to measure the Marconi's phase noise.
Today we just did a quick measurement but today next week we are going to repeat it more carefully.
Attached is a plot that shows the measurement calibrated for a UGF at about 60 Hz. The noise is compared to that specified by Wenzel for their crystal.
The noise is bigger than that of the MArconi alone locked to the Rubidium standard (see elog entry). We don't know the reason for sure yet.
We'll get back to this problem next week. |
Attachment 1: FSScrystalPhaseNoiseHigherGain.pdf
|
|
3284
|
Sat Jul 24 13:13:41 2010 |
rana, steve, alberto | Update | General | Initial Crane Inspection reveals flaws: wiring and oil |
The guy from KroneCrane (sp?) came today and started the crane inspection on the X End Crane. There were issues with our crane so he's going to resume on Monday. We turned off the MOPA fur the duration of the inspection.
- None of our cranes have oil in the gearbox and it seems that they never did since they have never been maintained. Sloppy installation job. The crane oiling guy is going to come in on Monday.
- They tried to test the X-End crane with 2500 lbs. (its a 1 ton crane). This tripped the thermal overload on the crane as intended with this test. Unfortunately, the thermal overload switch disabled the 'goes down' circuit instead of the 'goes up' circuit as it should. We double checked the wiring diagram to confirm our hypothesis. Seems the X-End crane was wired up incorrectly in the first place 16 years ago. We'll have to get this fixed.
The plan is that they will bring enough weight to test it at slightly over the rating (1 Ton + 10 %) and we'll retry the certification after the oiling on Monday. |
3283
|
Fri Jul 23 21:35:48 2010 |
rana | Update | CDS | Rerouted RFM around c1lsc, took RFM card out of c1lsc |
I just realized that an unfortunate casualty of this LSC work was the deletion of the slow controls for the LSC which we still use (some sort of AUX processor). For example, the modulation
depth slider for the MC is now in an unknown state. |
3282
|
Fri Jul 23 21:14:29 2010 |
Rana | Update | PSL | Reference Cavity Insulation |
I wrapped another ~3 layers onto there. It occurs to me now that we could just get some 2mm thick copper plates made to fit over the stainless steel can.
They don't have to completely cover it, just mostly. I also took the copper circles that Steve had made and marked them with the correct beam height
and put them on Steve's desk. We need a 1" dia. hole cut into these on Monday.
To compensate for the cooling during my work, I've set the heater for max heating for 1 hour and then to engage the temperature servo.
I also noticed the HEPA VARIAC on the PSL was set to 100. Please set it back to 20 after completing your PSL work so that it doesn't disturb the RC temperature.. |
3281
|
Fri Jul 23 18:55:25 2010 |
kiwamu | Update | SUS | PRM and TT installed |
|
3280
|
Fri Jul 23 16:02:16 2010 |
Rana | Update | PSL | Reference Cavity Insulation |
This is the trend so far with the copper foil wrapping. According to Megan's calculation, we need ~1 mm of foil and the thickness of each layer is 0.002" (1/20th of a mm), so we need ~20 layers. We have ~5 layers so far.
As you can see the out-of-loop temperature sensor (RCTEMP) is much better than before. We need another week to tell how well the frequency is doing -
the recent spate of power cycles / reboots of the PSL have interrupted the trend smoothness so far. |
Attachment 1: Untitled.png
|
|
3279
|
Fri Jul 23 16:00:35 2010 |
steve | Update | SAFETY | crane load test tomorrow |
All 3 cranes will be load tested at 1 ton tomorrow morning between 9am and 2pm
Do not come to the 40m lab during this period. We may disturb your experiment.
Please prepare your touchy set ups to take this test. |
3278
|
Fri Jul 23 15:54:38 2010 |
kiwamu | Update | SUS | PRM and TT installed |
[ Jenne, Koji and Kiwamu]
We have installed the PRM and the tip-tilt (TT) in the BS chamber.
We have started the in-vac work which takes about a week.
Today's mission was dedicated to installing the PRM and two TTs, one for the PRC and the other for the SRC, on the BS table in the chamber.
The work has been smoothly performed and we succeeded in installation of the PRM and a TT for the PRC.
But unfortunately the other TT got broken during its transportation from Bob's clean room.
(what we did)
(1) opened the light door of the BS chamber.
(2) moved the BS tower to the right position according to Koji's layout drawing.
- Prior to this work we screwed down the earthquake stops so that the mirror is fixed to the tower. Also we disabled the watchdog.
- When moving it we used an allen key as a lever with an screw as a fulcrum. This idea was suggested by Jenne and it really worked well.
The reason why we used this technique is that if we slide the tower by hands the tower can't go smoothly and it may sometimes skips.
After that we checked the postion from some reference screw holes by using a caliper and we made sure that it was on the right position.
(3) removed all of the square-shaped mirrors.
- After this removal the mirrors were wrapped by aluminum foils and put in a usual clear box.
(4) removed some optics because they had made the chamber space crowded.
- These were also wrapped by aluminum foils and put in the box. Later we will put them back to the BS table.
(5) brought the PRM tower from the Bob's clean room and put it on the BS table.
- The position of the PRM were coarsely aligned since we still don't have any 1064 beam going through the PRM.
(6) brought two TTs also from Bob's clean room and put one of the TTs on the table.
- The position of the installed TT was coarsely adjusted.
- After we brought them we removed the aluminum foils covering the TTs and we found the wire of a TT got broken.
It may have been damaged during its transportation from Bob's room because it was fine before the transportation.
(7) closed the door
(the next things to do)
* Installation of the OSEMs to the PRM
* Installation of the pick off mirror and its associated optics
* Arrangement of the pzt mirror
|
Attachment 1: DSC_2275.JPG
|
|
3277
|
Fri Jul 23 15:32:02 2010 |
steve | Update | PEM | control room AC set point changed |
The control room temp is warmer than usual. The heat exchanger Office Pro 18 set point was lowered from 70 to 68F yesterday.
The MOPA headtemp is higher also. The Neslab chiller bath temp peaks around 21.6 C daily. This should be rock solid 20.00 C
It did not have any effect.
Now, I have just lowered the thermostat setting of room 101 from 73 to71F I hope Koji can take this. |
3276
|
Fri Jul 23 14:26:01 2010 |
Gopal | Update | Optic Stacks | Simple Frequency Response Measurements in COMSOL |
Over the past couple days, I discovered a simple, direct method for calculating frequency responses with a combination of COMSOL and any plotter such as Excel or MatLab. The simple case of rectangular prism of steel was analyzed using this method; details will be posted shortly on the COMSOL Wiki page. The frequency response matched theoretical reasoning: the bar acts as a simple mechanical low-pass filter, rapidly attenuating driving frequencies at the base beyond the first eigenmode.
It therefore shouldn't be too difficult to extend this analysis to the MC1/MC3 stack. The many eigenfrequencies will produce a more complicated transfer function, and so more data points will be taken.
The major shortcoming of this method involves dealing with the imaginary components of the eigenfrequencies. As of now, I haven't found a way of measuring the phase lag between the drive and the response. I also haven't found a way of changing the damping constants and therefore playing with phase components.
|
3275
|
Fri Jul 23 12:33:51 2010 |
josephb | Update | CDS | megatron, c1iscex, c1sus firewalled |
Kiwamu and I strung an ethernet cable from the new 1X7 rack to the 1X3 rack. The cable is labeled c1iscex-daq on both ends. This cable will eventually connect c1iscex's second ethernet port to the daq router. However, for today, it plugged into the primary ethernet port and is going to a linksys router. This is the same linksys router we used to firewall megatron.
The idea is to place megatron, c1sus, and c1iscex behind the firewall to prevent any problems with the currently running while doing RFM nework tests.
The way to get into the firewalled sub-network is to ssh into megatron. The router will forward the ssh to megatron. Inside the network, the computers will have the following IPs. Router is 192.168.1.1, megatron is 192.168.1.2, c1sus is 192.168.1.3, and c1iscex is 192.168.1.4.
|
3274
|
Fri Jul 23 10:16:44 2010 |
josephb | Update | CDS | Rerouted RFM around c1lsc, took RFM card out of c1lsc |
I re-routed around the c1lsc machine this morning. I turned the crate off, and disconnected the transmission fiber from c1lsc (which went to the receiver on c1asc). I then took the receiving fiber from c1lsc and plugged it into the receiver on c1asc.
I pulled out the c1lsc computer from the VME crate and pulled out the RFM card, which I needed for the CDS upgrade. I then replaced the lsc card back in the crate and turned it back on. Since there hasn't been a working version of the LSC code on linux1 since I overwrote it with the new CDS lsc code, this shouldn't have any significant impact on the interferometer.
I've confirmed that the RFM network seems to be in a good state (the only red lights on the RFM timing and status medm screen are LSC, ASC, and ETMX). Fast channels can still be seen with dataviewer and fb40m appears to still be happy.
The RFM card has found its new home in the SUS IO Chassis. The short fiber that used to go between c1asc and c1lsc is now on the top shelf of the new 1X3 rack. |
3273
|
Fri Jul 23 08:18:03 2010 |
josephb | Update | CDS | Not enough room in IO chassis for RFM card - need to swap PMC to PCIe adapters |
The End IO chassis have small trenton boards, which apparently only have 5 usuable PCI slots, even though there are 6 on the board. This is because of the way the the host interface board is setup and its closeness to the 2nd to last PCI slot
The PMC to PCIe adapters I was handed by Jay for use with the RFM cards require a 4 pin power connection at the top, which are not available inside the thin 1U computers.
The only solution I can come up with is swap the PMC to PCIe adapters for the RFM cards with adapters for some of the already installed ADCs and DACs which do not require power directly from the power supply. This should make it possible to mount the RFM card in the computer, at least for the ends. Since the SUS and IOO chassis will have more slots available than needed, the RFM cards can be slotted into those. The SUS has to fit in the chassis since the computer will have the Infiband host adapter and a dolphin connector for talking to the LSC machine.
There is still the problem of actually getting the RFM card into the computer, but that should be possible with a little bit of bending of the left side of the computer frame. |
3272
|
Fri Jul 23 08:15:59 2010 |
steve | Update | PSL | ref cavity ion pump |
The ref cavity ion pump was running at 7.7kV instead of 5kV
This Digitel SPC-1 20 l/s ion pump should be running at 5kV |
3271
|
Fri Jul 23 00:13:11 2010 |
rana | Update | PSL | Problem NOT REALLY Solved |
So...who was working around the PSL rack this morning and afternoon? Looks like there was some VCO phase noise work at the bottom of
the rack as well as some disconnecting of the Guralp cables from that rack. Who did which when and who needs to be punished? |
3270
|
Thu Jul 22 18:18:54 2010 |
Alberto | Update | PSL | Problem Solved |
Quote: |
Quote: |
It looks like something wrong happened around the PSL front end. One of the PSL channel, C1:PSL-PMC_LOCALC, got crazy.
We found it by the donkey alarm 10 minutes ago.
The attached picture is a screen shot of the PMC medm screen.
The value of C1:PSL-PMC_LOCALC ( middle left on the picture ) shows wired characters. It returns "nan" when we do ezcaread.
Joe went to the rack and powered off / on the crate, but it still remains the same. It might be an analog issue (?)
|
The problem seems to be a software one.
In any case, Kiwamu and I looked at the at the PMC crystal board and demod board, in search of a possible bad connection. We found a weak connection of the RG cable going into the PD input of the demod board. The cable was bent and almost broken.
I replaced the SMA connector of the cable with a new one that I soldered in situ. Then I made sure that the connection was good and didn't have any short due to the soldering.
|
[Alberto, Koji]
By looking at the reference pictures of the rack in the wiki, it turned out that the Sorensen which provides the 10V to the 1Y1 rack was on halt (red light on). It had been like that since 1.30pm today. It might have probably got disabled by a short somewhere or inadvertently by someone working nearby it.
Turning it off and on reset it. The crazy LO calibrated amplitude on the PMC screen got fixed.
Then it was again possible to lock PMC and FSS.
We also had to burtrestore the PSL computer becasue of the several reboots done on it today. |
3269
|
Thu Jul 22 15:59:29 2010 |
Alberto | Update | PSL | PSL front end machine |
Quote: |
It looks like something wrong happened around the PSL front end. One of the PSL channel, C1:PSL-PMC_LOCALC, got crazy.
We found it by the donkey alarm 10 minutes ago.
The attached picture is a screen shot of the PMC medm screen.
The value of C1:PSL-PMC_LOCALC ( middle left on the picture ) shows wired characters. It returns "nan" when we do ezcaread.
Joe went to the rack and powered off / on the crate, but it still remains the same. It might be an analog issue (?)
|
The problem seems to be a software one.
In any case, Kiwamu and I looked at the at the PMC crystal board and demod board, in search of a possible bad connection. We found a weak connection of the RG cable going into the PD input of the demod board. The cable was bent and almost broken.
I replaced the SMA connector of the cable with a new one that I soldered in situ. Then I made sure that the connection was good and didn't have any short due to the soldering. |
3268
|
Thu Jul 22 14:07:20 2010 |
kiwamu | Update | PSL | PSL front end machine |
It looks like something wrong happened around the PSL front end. One of the PSL channel, C1:PSL-PMC_LOCALC, got crazy.
We found it by the donkey alarm 10 minutes ago.
The attached picture is a screen shot of the PMC medm screen.
The value of C1:PSL-PMC_LOCALC ( middle left on the picture ) shows wired characters. It returns "nan" when we do ezcaread.
Joe went to the rack and powered off / on the crate, but it still remains the same. It might be an analog issue (?) |
Attachment 1: PSL-PMC2010-07-22.png
|
|
3267
|
Thu Jul 22 13:44:47 2010 |
Jenna | Update | PEM | Guralp seismometer |
One of the Guralps [Gur2] has been taken to the atf gyro lab, along with the breakout box.
Edit by Jenne: This means that we have no working seismometers in the 40m lab right now, so don't worry if you're looking for seismo data and you can't find any. The 6 accelerometers should all still be up and running. |
3266
|
Thu Jul 22 08:07:27 2010 |
steve | Update | VAC | vent finished |
Quote: |
The vent has been finished successfully in this morning.
The vent was finished successfully this morning.
|
Thanks to Kiwamu, Alberto and Koji |
Attachment 1: 30hvent.jpg
|
|
3265
|
Thu Jul 22 07:19:56 2010 |
Alberto | Update | Treasure | Monsters, LNVR, and Phase noise |
"They (shellfish) shall be an abomination to you; you shall not eat their flesh, but you shall regard their carcasses as an abomination." (Leviticus 11:11) |
3264
|
Thu Jul 22 03:08:27 2010 |
Koji | Update | IOO | Vent and MC lock |
Summary
- The vacuum chambers have been vented.
- The north heavy door of the BS chamber has been opened by Genie (not by the crane).
It was replaced by the light door. The door is currently closed.
- The MC has been locked with 20mW incident and aligned. MC REFL was left unchanged but lock was able to be achieved.
- The optics before MCT CCD and MCT QPD have been adjusted for the low power operation.
Details
- The first HWP for the variable optical attanuator (HWP/PBS/HWP pair) was set to be 86deg from the maximum transmission at 126deg.
The incident power of 19mW has been measured.
- The PSL mechanical shutter has been manually opened. Two other beam blocks has been removed.
- I found the MC was totaly misaligned with no resonance.- I tried to align it based on the previous OSEM values but in vain.
• How to align the MC mirrors from the scratch
- MC1 has been aligned so as to maximize REFL PD and DC signal of WFS QPDs.
- MC3 has been aligned by looking at the scattered light on the MC2 frames. The spot is centered on the MC2 approximately.
- MC2 has been aligned so that any resonance is seen in MC_F.
• Modification of the MCT optics
- The ND filter before the MCT was removed.
- The Y1-45S mirror before the MCT CCD, which is also used to steer the beam to the MCT QPD path, was replaced to BS1-50-45P.
The reason I used 45P is to obtain higher reflectivity. Because S has higher reflectivity than P in the each layer, I expected to have higher reflectivity for S than 50%.
- The MC REFL path has not been untouched.
• Modification of the servo
- The lock was attempted after alignment of the mirrors. Here how to lock the MC is described below.
1. Run script/MC/mcloopson
2. Open the MC Servo screen in MEDM
3. Change the input gain from 6dB to 22dB.
Change offset from 0.78 to -0.464 (such that the length output has no offset).
Change VCO Gain from 3dB to 21dB
Change MC Length path Gain from 0.3 to 1.6
|
3263
|
Thu Jul 22 01:02:08 2010 |
rana | Update | Treasure | Monsters, LNVR, and Phase noise |
On Picasa |
3262
|
Wed Jul 21 19:11:18 2010 |
Dmass | Update | Green Locking | locked |
What did you use to filter the 2f components from your error signal? A homemade low pass or what?
Kiwamu:
I am using a homemade low pass filter.
It's just a RC passive LPF with the input impedance of 50 Ohm. |
3261
|
Wed Jul 21 17:41:17 2010 |
Gopal | Configuration | Optic Stacks | Pictures of Stacks |
Now that venting is complete, this is a request for anyone who opens any chamber:
1) Please notify me immediately so I can take pictures of the stacks in that chamber.
2) If I'm not around, please take a few pictures for me. I'm most interested in the shape, number of layers, size, and damper arrangements of each stack.
This is most important for the MC1/MC3 chamber, MC2 chamber, and BS/ITMX/ITMY chambers.
Thanks! |
3260
|
Wed Jul 21 15:43:38 2010 |
Megan | Summary | PSL | Copper Layer Thickness on the Reference Cavity |
Using the equation for thermal resistance
Rthermal = L/(k*A)
where k is the thermal conductivity of a material, L is the length, and A is the surface area through which the heat passes, I could find the thermal resistance of the copper and stainless steel on the reference cavity. To reduce temperature gradients across the vacuum chamber, the thermal resistance of the copper must be the same or less than that of the stainless steel. Since the copper is directly on top of the stainless steel, the length and width will be the same for both, just the thickness will be different (for ease of calculation, I assumed flat, rectangular strips of the metal). Assuming we wish to have a thermal resistance of the copper n times less than that of the stainless steel, we have
RCu = RSS/n
or
L/(kCu*w*tCu) = L/(kSS*w*tSS*n)
so that
tCu/tSS = n*kSS/kCu
We know that kSS = 401 W/m*K and KCu = 16 W/m*K, so
tCu/tSS = 0.0399*n
By using the drawings for the short reference cavity vacuum chamber (the only one I could find drawings for online) I found a thickness of the walls of 0.12 in or 0.3048 cm. So for the same thermal resistance in both metals, the copper must be 0.0122 cm thick and for a thermal resistance 10 times less, it must be 0.122 cm thick. So we will have to keep wrapping the copper on the vacuum chamber! |
3259
|
Wed Jul 21 12:41:31 2010 |
kiwamu | Update | VAC | vent finished |
The vent has been finished successfully in this morning.
The vent was finished successfully this morning. |
3258
|
Wed Jul 21 12:20:58 2010 |
Razib | Update | Phase Camera | Weekly update |
This past week I have worked on the following:
1. Setting up the infrastructure to do noise analysis: We added a temporary channel on the DAQ to connect to the PD 55 which we are using to take the power measurement. Before that, I connected the PD55 to an oscilloscope and recorded the power.

The power at PD55 as measured using the oscilloscope = 600 µV.
Then I tried to calibrate the channel by sending up a signal from the function generator and measuring up the offset.. However, the channels seems noisy enough, especially due to electronics noise as suggested by the measurements and FFT calculation.
2. I worked on trying to sync the data acquisition of the PD and the CAM. After sometime spent on fiddling with the software method such as taking images at stamped time and then lining them up with the daq timestamps, I found a hardware method as suggested by Aidan. It was putting up a shutter (Uniblitz shutter and driver VMMD1) in the setup. I synced the shutter with the camera for which I had to tear apart the previously made trigger box and add a sync output from the camera (took a while as I also had to make a new cable).
3. I worked (still working) on making a differential amplifier to blow up the signal from the PD.
|
3257
|
Wed Jul 21 12:20:29 2010 |
josephb, kiwamu | Update | Computers | Megatron temporarily disconnected, c1iscex firewalled, green FE test |
We are moving towards a first test of getting Kiwamu's green locking signals into the new front end at the new X end, as well as sending signal out to the green laser temperature control.
Towards that end, we borrowed the router which we were using as a firewall for megatron. At the moment, megatron is not connected to the network. The router (a linksys N wire router), was moved to the new X end, and setup to act as a firewall for the c1iscex machine.
At this point, we need to figure which channels of the DAC correspond to which outputs of the anti-imaging board (D000186) and coil driver outputs. Ideally, we'd like to simply take a spare output from that board and bring it to the laser temperature control. The watchdogs will be disabled when testing to avoid any unfortunate mis-sent signals to the coils. It looks like it should be something like channels 6,7,8 are free, although I'm not positive if thats the correct mapping or if there's a n*8 + 6,7,8 mapping.
The ADC should be much easier to determine, since we only have a single 16 channel set coming from the lemo breakout box. Once we've determined channels, we should be all set to do a test with the green system. |
3256
|
Wed Jul 21 12:03:14 2010 |
Katharine, Sharmila | Update | WIKI-40M Update | weekly update |
This past week, we levitated our small cylindrical magnet (with the flag made from heat shrink). Though the levitated magnet didn't appear very jittery to the eye, we looked at the PD current on the scope and could see oscillations that corresponded to the flag hitting the sides of the OSEM. The oscillations were more pronounced as we gently hit/vibrated the lab bench, and by pounding on the bench Rana knocked the levitated magnet completely out of the setup. So, we're currently working on building a new, stabler mount. The biggest challenge here is fixing the OSEM in place, but we're experimenting with different optics pieces to see which is the most stable for our purposes. Jenne taught us how to make through holes using the drill press so we can add slats of aluminum to adjust the height of the OSEM mount. We also plan to fix some heavy plates to the bottom of our system to decrease its vibration frequency.
We also calculated the transfer function of our circuit, which seems to match the measured frequency response to within a small factor. We're playing with Rana's Simulink models and are currently modeling our own system to determine what gains we would expect to use and get a better understanding of our circuit.
Once our system is successfully mounted stably, we plan to experimentally observe the effects of changing the gain and integrator by looking at time series measurements of the PD current, as well as using the spectrum analyzer to compare the frequency response of our system at different gain settings.
|
3255
|
Wed Jul 21 11:57:59 2010 |
Gopal | Update | WIKI-40M Update | 7.14.10-7.21.10 Weekly Update |
Summary of this week's activities:
7/14: Analytical calculation of Viton spring constant; updated Viton values in models; experimental confirmation of COMSOL eigenfrequencies (single stack layer)
7/15: Extensions to 2-, 3-, and 4-layer stack legs. Eigenfrequency characterizations performed for each level. Meshing issues with 4-layer stack prevented completion.
7/19: Debugged the 4-layer stack. Turned out to be a boundary condition issue because of non-sequential work-plane definitions. Successful characterization of single-leg eigenfrequencies.
7/20: Prototype three-legged stack completed, but dimensions are incorrect. Read Sievers paper for details of triple-legged stack. Sorted through many stack design binders in efforts to distinguish IOC/OOC, BSC/ITMX/ITMY, MC1/MC3, and MC2 dimensions.
7/21: Researched frequency domain analysis testing in COMSOL. Attempting to first find transfer function of a single-layer stack --> currently running into some run-time errors that will need some more debugging in the afternoon. |
3254
|
Tue Jul 20 23:52:36 2010 |
kiwamu | Update | VAC | update: slow vent has started |
The vent is still going on. At this moment the pressure inside of the chambers is about 630 Torr.
Koji and I have replaced the 2nd instrument grade compressed air cylinder by the 3rd cylinder around 9 pm.
So far the vent speed has been nicely kept at about 1 Torr / min. |
Attachment 1: untitled.png
|
|
3253
|
Tue Jul 20 18:29:43 2010 |
nancy | Update | IOO | QPD installed behind the MC2 |
Yesterday I installed teh QPD on the table behind MC2, and observed teh signal on it.
The MC_leak is directed to it by a steering mirror.
I used the A2L_MC2 script to minimise teh pitch and yaw gains, and estimated teh spot position on teh MC2 using that.
This spot position was aligned to the center of teh QPD.
In the night while before taking measurements, I decided to turn off the Wavefront Sensor Servos, but just after that, the MC alignment went very bad, and I could not align it in the next 2 hours.
For some reason, the MC was really mad the whole day yesterday, and was getting misaligned again and again, even when the WFS feedback was on.
The table also had another IR laser in it, which I and Koji switched off.
I will continue measuring once we pump down again.
For now, I am analysing teh QPD circuit Transfer Function. |
3252
|
Tue Jul 20 17:38:16 2010 |
Gopal | Configuration | Optic Stacks | Stack Type Clarifications |
Some clarification is warranted regarding the different shapes of stacks. Corrections are appreciated:
1) The single-leg stack that I just completed should function as a working model for the IO, OO, and MC1/3. Rana commented, however, that the current dimensions are slightly off for MC1/3 (which makes sense since I could only find drawings for the IOC). If anyone knows the whereabouts of similar drawings for MC1/3, I'd much appreciate it.
2) A triple-leg stack can model the BS, ITMX, and ITMY chambers. I don't have exact dimensions for these, but I can make decent approximations from to-scale stack drawings. I'll probably work on a model for this style next, since at least I have some information regarding this version.
3) The MC2 chamber has its own stack model, about which I haven't found any drawings in the binders. I can't start on MC2C at all until I find such drawings. |
3251
|
Tue Jul 20 15:06:48 2010 |
josephb, bob | Omnistructure | General | New speakers in ceiling in control room |
After the crane training, Bob attached speakers to the ceiling right next to the projector, for use with presentations. |
3250
|
Tue Jul 20 11:55:15 2010 |
Sharmila,Katharine | Update | elog | Maglev |
We plotted the transfer functions for the maglev control circuit and compared them with the plots from the spectrum
analyzer. We were stuck for sometime because
1) we had wrongly entered the value of one of the resistors which was off by a factor of 2000.
2) The plots were not done in right units. So we couldn't figure out differences quite well.
The two plots are shown below. We are still off by a factor of 3 which we'll figure out soon.

|
3249
|
Tue Jul 20 11:49:31 2010 |
Jenne | Update | SUS | Some Suspensions not damping |
[Jenne, Koji]
I moseyed into the control room this morning, to find ITMX and ITMY both with their watchdogs tripped. ITMY (new convention) wouldn't damp. Koji discovered that there was a sign flip in 2 of the sensors. A set of reboots of c1susvme1&2 fixed the problem.
A side note: For the ETMs, the OSEM sensor readouts are gigantic (~20,000), whereas for the similar channels on all other optics, the readouts are on the order of 1. After some looking around, it seems that this is just the way things have been (for at least 100 days), and the filters in the SUSPOS and other SUS filter banks have a high pass filter to take care of this. It's weird, but it seems to be the way it is, and the ETMs damp, so it's all good. |
3248
|
Tue Jul 20 08:04:31 2010 |
steve | Configuration | VAC | slow vent has started |
The PSL shutter was closed. The beam path blocked two places. High voltage power supplies to IOO and OMC PZT were checked to be off. Oplevs are off.
The south arm green cavity was misaligned yesterday
We would like to keep the vent speed at 1 torr / min. I'm venting with N2 now up to 25 PSI. We have 3 cylinder of instrument grade air inside the lab. Additional supply will arrive later. It can be as late as 1pm |
3247
|
Mon Jul 19 21:47:36 2010 |
rana | Summary | DAQ | DAQ timing test |
Since we now have a good measurement of the phase noise of the Rb clock Marconi locked to the Rb clock, I wanted to use that to check out the old DAQ system:
I used Megan's phase noise setup - Marconi #2 is putting out 11000013 Hz at 13 dBm into the ZP-3MH mixer. Marconi #1 is putting out 3 dBm at 11000000 Hz into the RF input.
The output goes through a 50 Ohm load and then a Mini-Circuits BNC LP filter (either 2 or 5 MHz). Then an SR560 set for low noise, G = 5, AC coupling, 1-pole LP @ 1 kHz.
This SR560 output goes into the channel C1:IOO-MC_DRUM1 (which is sampled at 16384 Hz with ICS-110B after the usual Sander Liu AA chassis containing the INA134s). |
3246
|
Mon Jul 19 16:11:17 2010 |
Gopal | Update | Optic Stacks | Eigenfrequency Analysis of Full Stack |
Expanding on the single-layer model, I added the second, third, and fourth layers to the stack in COMSOL. Eigenfrequency analysis run times increased exponentially as the model multiplied in complexity. The following images document the some of the important eigenfrequencies:
First Eigenmode: y-translational, 3.34 Hz:

Second Eigenmode: x-translational, 3.39 Hz:

Third Eigenmode: z-rotational, 3.88 Hz:

Sixth Eigenmode: z-translational, 8.55 Hz:

As expected, the eigenfrequencies are generally lower, but still in the same range, as the single-layer model, because of greater mass but constant weight-per-spring distribution.
Next Steps:
1) Extend a single stack to the full stack system, which consists of three stacks like this. Perform similar eigenmode analysis.
2) Analyze the mirror suspension system and incorporate a similar pendulum on the top plate.
3) Make transfer function measurements between seismic and mirror motions. |
3245
|
Mon Jul 19 14:16:01 2010 |
nancy, koji | Update | IOO | QPD Response Transfer Function |
Quote: |
Friday night myself and Koji measured the Transfer function of the QPD circuit at MC2 side using a chopper . Following was our procedure :
We connected some wires at the input and output of teh filter circuit to one of the segment of teh QPD. - seg 2.
A laser light was shined on to the QPD, it was pulsed using a chopper. The frequency of rotation of teh chopper was varied.
These wires were then fed to the spectum analyser , and a transfer funstion was observed, It was nearly a low pass filter
The chopper frequency was then made variable by giving the chopper a signal from teh spectrum analyser. This signal just swiped a large range of the rpm of the chopper.
Now the input signal looked like a sine wave of varying frequency. the transfer functino looked like a perfect LPF, with a small SNR.
Attaching the plot of the TF in the next e-log (this one is on windows and cant access /cvs/cds)
|
 |